OpenStudy (anonymous):

if X= y+3y^2+4y^3, then y^1=

OpenStudy (anonymous):

Not getting your question... Sorry..

OpenStudy (anonymous):

No idea ......what do you wanna solve ..

OpenStudy (anonymous):

differentiation

OpenStudy (anonymous):

So : \[X = y+ 3y^2 + 4y^3\] We are to find : \(y'\) Right ??

OpenStudy (anonymous):

right

OpenStudy (anonymous):

y"=dy/dx= 1/ 12y^2+6y+1